shift sum 10/2/99

Theodoros Alexopoulos (theoalex@fnal.gov)
Sat, 2 Oct 1999 15:42:21 -0500 (CDT)

Shift Summary for Day Shift, 10/2/99

On Shift: Mark Timko, Ashkan for the first 4 hours, Theo for the rest of the shift

Hours of beam available at all intensities: 7.5h
Hours of beam available at nominal intensity: 7.5h
Hours of E799 data (special runs not included): 7.0h
Hours of special runs (laser scans, muon runs...): 0h

Accounting of lost time: tape change, kfrends

========================================================================
Runs ended this shift:
Run 14793 - BEAM_799 - 23M events
Run 14794 - BEAM_799 - in progress

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:
smooth E799 datataking (SEM=8.5E12). A few spills with dead time,
and ktevana errors; kfrend fixes everything...

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14793 7:11 ktevana errors; kfrend it.

7:15 Mount UPX004 for online split.

7:11 ktevana errors; kfrend it.

12:50 run ended

13:05 Install new DC T0 constants after Ashkan's request

14794 13:15 Start run; ktevana errors; kfrend it.

13:25 ktevana errors; kfrend it.

13:45 ktevana errors; kfrend it.

14:15 SEM is wondering around 8.0E12. Call MCR to bring it
back to the request level (8.5E12)

14:20 ktevana errors; kfrend it.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

BobT would like to be paged if there is no beam for >= 1 hour
(day/evening).

theo for Mark and Ashkan